how can we express electric field on conducting sphere.
1
Expert's answer
2020-09-21T08:33:02-0400

Let the charge on the sphere be Q and the radius of the sphere be R.


We may apply the Gauss's law (see https://en.wikipedia.org/wiki/Gauss%27s_law) to the sphere. The total flux "\Phi" through the surface of the sphere ("4\pi R^2" ) is proportional to the total charge inside the sphere:

"\Phi = E \cdot S = E \cdot 4\pi R^2 = \dfrac{Q}{\varepsilon_0}" . Therefore, the field of the sphere will be

"E = \dfrac{Q}{4\pi\varepsilon_0 R^2}," where "\varepsilon_0" is the vacuum permittivity (ε0 = 8.8541878128(13)×10−12 F⋅m−1).


More detailed derivation can be found in farside.ph.utexas.edu/teaching/302l/lectures/node25.html for the negative charge.